2. (a) Prove that K = (Z/3Z)[x]/(x3 + 2x + 1) is a field. (b) Find all the roots of the polynomial f(t) = + + 2t +1 in K.
Answer #1

@ @ Let PC) = x2+2x+1 e 2 [] & deg (P(x)) = 3 where Z = 72 = {0, 1, 2} Now Plo) = 0 +2.0+1=1+0 in 2 POD)- +2.1+ 1 =1+2+1=4 =For t=14& set f (148)= (1+2 3+2 (1+x) + 1 = 1 +32 +3x2+x3 + 2 + 2x + 1 = 1+x3+2 +26+1 -132010 Z) = 23+2x+1+2 = x²+2x+1 - (:32